OpenStudy (anonymous):

2x + 1 greater than or equal to 5 or -3x > -9 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (texaschic101):

2x + 1 > = 5 2x >= 5 - 1 2x > = 4 x > = 2 -3x >= -9 x <= 3 x is greater then or equal to 2 and x is less then or equal 3 line graph : closed circle on 2, closed circle on 3, shading in between.